# Evolution of AI Agent Registry Solutions: Centralized, Enterprise, and Distributed Approaches

### v2

@@ 1 Introduction
Traditional mechanisms like DNS and static service catalogs are poorly suited to agent ecosystems, which demand dynamic discovery, verifiable metadata, and privacy-preserving interactions. Legacy systems assume fixed endpoints and ownership-based trust models, lacking the flexibility and cryptographic assurances needed for agents that rotate capabilities, change locations, and form ephemeral collaborations.
@@
2 Background
@@
This shift introduces significant challenges for discovery and coordination. The current Internet stack built on DNS, IP addressing, and certificate authorities was not designed to handle trillions of fast-moving, self-directed agents. Limitations in revocation latency, state propagation, identity verification, and routing scale all become critical bottlenecks.
@@
2.1 Design Evaluation Dimensions
@@ 3.4 Maintenance
4 Agent2Agent (A2A) Protocol
@@ 4.4 Maintainability
5 AGNTCY Agent Directory Service
@@
Building on the four layers above, ADS introduces a semantic enrichment path. Hierarchical skill taxonomies accelerate capability-scoped queries before DHT resolution. Replica metadata (latency, region, freshness) feeds a composite scoring function that produces deterministic, low-cost selection decisions. Location-independent OCI digests (sha256:<digest>) enable transparent multi-registry reconciliation: any registry serving the same digest becomes an interchangeable source. Digest equality, reinforced by Sigstore provenance attestations, enforces immutability and prevents substitution. Dynamic selection then avoids centralized choke points by coupling semantic relevance with Kademlia routing proximity.
@@ 5.2 Authentication
Server Authentication. This ensures trust among server peers and restricts unauthorized access. Core discovery can be left open (content trust is intrinsic), but operators often restrict which peers a directory node will accept or exchange routing/index data with. “Secure mode” introduces: (1) authenticated peer admission (static allowlist or dynamic attestation) before Kademlia session establishment; (2) mutual TLS with short-lived SPIFFE/SPIRE SVIDs or equivalent rotating credentials. These measures shrink the unsolicited attack surface (e.g., DDoS amplification/state exhaustion) and satisfy enterprise policies that forbid arbitrary federation.
5.3 Scalability
@@ 5.4 Maintainability
6 Microsoft Entra Agent ID
@@ 7 NANDA Index
    Dynamic Resolution Layer: Interprets metadata to route queries through static, rotating, or adaptive endpoints. This layer supports geo-aware routing, load balancing, and short-lived endpoint rotation (TTL 5–15 minutes) to meet performance, privacy, and resiliency requirements at scale.
@@ 7.4 Maintainability
Table 1: Comparison of Five Agent Registry Approaches
@@
Table 2: Agent Advertisement Models: Card vs. Facts
@@
Revocation Re-fetch interval VC-Status (<1s)
@@ 8 Comparative Evaluation
  • Table 1 compares registry solutions from MCP, A2A (Google), AGNTCY, Entra Agent ID and NANDA Index, analyzing their purpose, discovery paths, trust primitives, privacy mechanisms, endpoint freshness strategies, schema complexity, and best-fit use cases.
  • Table 2 provides a detailed feature-level comparison between Agent Cards (used in A2A) and Agent Facts (used in NANDA Index), highlighting key differences—including but not limited to metadata structure, endpoint modeling, cryptographic guarantees, and extensibility.
9 Evolution of Agent Registry Architectures
Agent registry systems have evolved from simple, file-based descriptions to distributed, cryptographically-verifiable registries with structured discovery protocols. This section outlines the progression across thtree key phases, each adding layers of interoperability, scalability, trust, and governance. Understanding these phases helps clarify both the design motivations behind existing systems and the trade-offs made at each maturity level.
9.1 Static, Isolated Discovery
@@
Examples: Google A2A’s /.well-known/agent.json,
9.2 Dynamic RESTfull APIs
@@
9.3 Verifiable Metadata and Federated Trust and AI agent quilt
@@ 10 Discussion
    Architectural Trade-offs Are Protocol-Specific. The "right" registry architecture depends heavily on deployment context. Enterprise environments with existing Azure AD infrastructure benefit from Entra Agent ID’s seamless integration and zero-maintenance approach. Open research communities and decentralized applications require the cryptographic guarantees and federated governance of NANDA Index. Protocol-specific ecosystems like MCP benefit from purpose-built registries that leverage existing authentication systems (GitHub OAuth) while maintaining simplicity.
@@
    Security Must Be Built-in, Not Bolted-on. All examined registries recognize that cryptographic integrity is foundational whether through W3C Verifiable Credentials (NANDA), DNS-TXT verification (MCP), or Azure AD’s enterprise security controls (Entra). However, AGNTCY ADS stands out with its Sigstore-backed keyless signing and optional DID-based identity, ensuring both data authenticity and server authentication. This approach prevents registry poisoning and enables privacy-preserving discovery patterns.
@@
    Interoperability Remains the Critical Gap. Despite their architectural differences, these registries serve overlapping use cases and will inevitably need to interoperate as agent ecosystems mature. Cross-protocol discovery, unified namespace management, and portable agent identities represent the next frontier for registry infrastructure development.
@@
11 The Path Ahead: A Switchboard for the Agentic Web
@@
To move beyond research, our teams are deploying a working solution in two phases, inviting researchers and practitioners to join: Phase 1 establishes the NANDA Index at MIT as a neutral, consortium-governed switchboard for cross-organizational agent coordination; Phase 2 adds semantic discovery and federation capabilities through AGNTCY’s Agent Directory, enabling interoperability with heterogeneous registry types via DHT-based routing and OCI-aligned content addressing.

### v3

@@ I Introduction
Traditional mechanisms like DNS and static service catalogs are poorly suited to agent ecosystems, which demand dynamic discovery, verifiable metadata, and privacy-preserving interactions [1]. Legacy systems assume fixed endpoints and ownership-based trust models, lacking the flexibility and cryptographic assurances needed for agents that rotate capabilities, change locations, and form ephemeral collaborations.
@@
In trust-critical domains such as healthcare, finance, and critical infrastructure, agent registries must answer: How do we verify an agent is who it claims to be? How do we ensure its advertised capabilities haven’t been tampered with? How do we revoke access when trust is violated? This paper evaluates how five registry approaches address these trust primitives.
II Background
@@
This shift introduces significant challenges for discovery and coordination. The current Internet stack built on DNS, IP addressing, and certificate authorities was not designed to handle trillions of fast-moving, self-directed agents. Limitations in revocation latency, state propagation, identity verification, and routing scale all become critical bottlenecks, particularly in trust-sensitive domains requiring zero-trust architectures and continuous verification [7, 8]. The fundamental question of whether to upgrade existing infrastructure or implement purpose-built agent registries represents a qualitative, not incremental, architectural transition [5].
@@
II-A Trust Requirements in Agent Registries
Agent registries must provide three pillars of trust:
  • Identity Assurance: Cryptographic binding between agent identifiers and their metadata, preventing impersonation and capability spoofing. Critical for healthcare agents handling protected health information (PHI).
  • Integrity Verification: Tamper-evident metadata and audit trails that detect unauthorized modifications [10]. Essential for regulatory compliance (HIPAA, FDA, GDPR).
  • Privacy Preservation: Discovery mechanisms that don’t leak sensitive information about capabilities, access patterns, or organizational relationships—particularly important in clinical research and cross-institutional collaborations.
These informed our evaluation framework in Section VIII.
II-B Design Evaluation Dimensions
@@ III-D Maintenance
IV Agent2Agent (A2A) Protocol
@@ IV-D Maintainability
V AGNTCY Agent Directory Service
@@
Building on the four layers above, ADS introduces a semantic enrichment path. Hierarchical skill taxonomies accelerate capability-scoped queries before DHT resolution. Replica metadata (latency, region, freshness) feeds a composite scoring function that produces deterministic, low-cost selection decisions. Location-independent OCI digests (sha256:<digest>) enable transparent multi-registry reconciliation: any registry serving the same digest becomes an interchangeable source. Digest equality, reinforced by Sigstore provenance attestations, enforces immutability and prevents substitution. Dynamic selection avoids centralized choke points by coupling semantic relevance with Kademlia routing proximity.
@@ V-B Authentication
Server Authentication. This ensures trust among server peers and restricts unauthorized access. Core discovery can be left open (content trust is intrinsic), but operators often restrict which peers a directory node will accept or exchange routing/index data with. “Secure mode” introduces: (1) authenticated peer admission (static allowlist or dynamic attestation) before Kademlia session establishment; (2) mutual TLS with short-lived SPIFFE/SPIRE SVIDs or equivalent rotating credentials. These measures shrink the unsolicited attack surface (e.g., DDoS amplification/state exhaustion) and satisfy enterprise policies that forbid arbitrary federation. Such zero-trust peer authentication patterns are increasingly required in regulated environments [8].
V-C Scalability
@@ V-D Maintainability
V-E Enterprise Deployment Considerations
The NANDA architecture has been explored in enterprise environments [15], demonstrating cross-protocol interoperability with MCP, A2A, and standard HTTPS communications. The framework proposes Zero Trust Agentic Access (ZTAA) principles, extending traditional Zero Trust Network Access to address autonomous agent security challenges across heterogeneous protocol environments. This deployment experience informs the design choices outlined above, particularly regarding federation, policy enforcement, and credential management at scale.
VI Microsoft Entra Agent ID
@@ VII NANDA Index
    Dynamic Resolution Layer: Interprets metadata to route queries through static, rotating, or adaptive endpoints [16]. This layer supports context-aware, real-time endpoint selection based on geographic location, system load, agent capabilities, and security threats. The AdaptiveResolver architecture enables short-lived endpoint rotation and negotiation of trust, quality of service, and resource constraints to meet performance, privacy, and resiliency requirements at scale.
@@ VII-D Maintainability
TABLE I: Comparison of Five Agent Registry Approaches
@@
TABLE II: Agent Advertisement Models: Card vs. Facts
@@
Revocation Re-fetch interval VC-Status (¡1s)
@@ VIII Comparative Evaluation
  • Table I compares registry solutions from MCP, A2A (Google), AGNTCY, Entra Agent ID and NANDA Index, analyzing their purpose, discovery paths, trust primitives, privacy mechanisms, endpoint freshness strategies, schema complexity, and best-fit use cases.
  • Table II provides a detailed feature-level comparison between Agent Cards (used in A2A) and Agent Facts (used in NANDA Index), highlighting key differences—including but not limited to metadata structure, endpoint modeling, cryptographic guarantees, and extensibility.
IX Evolution of Agent Registry Architectures
Agent registry systems have evolved from simple, file-based descriptions to distributed, cryptographically-verifiable registries with structured discovery protocols. This section outlines the progression across thtree key phases, each adding layers of interoperability, scalability, trust, and governance.
IX-A Static, Isolated Discovery
@@
Example: Google A2A /.well-known/agent.json
IX-B Dynamic RESTfull APIs
@@
IX-C Verifiable Metadata and Federated Trust and AI agent quilt
@@ X Discussion
    Architectural Trade-offs Are Protocol-Specific. The ”right” registry architecture depends heavily on deployment context. Enterprise environments with existing Azure AD infrastructure benefit from Entra Agent ID’s seamless integration and zero-maintenance approach. Open research communities and decentralized applications require the cryptographic guarantees and federated governance of NANDA Index. Protocol-specific ecosystems like MCP benefit from purpose-built registries that leverage existing authentication systems while maintaining simplicity.
@@
    Security Must Be Built-in, Not Bolted-on. All examined registries recognize that cryptographic integrity is foundational whether through W3C Verifiable Credentials (NANDA), DNS-TXT verification (MCP), or Azure AD’s enterprise security controls (Entra). However, AGNTCY ADS stands out with its Sigstore-backed keyless signing and optional DID-based identity, ensuring both data authenticity and server authentication. This approach prevents registry poisoning and enables private discovery patterns.
@@
    Interoperability Remains the Critical Gap. Despite architectural differences, these registries serve overlapping use cases and will inevitably need to interoperate as agent ecosystems mature. Cross-protocol discovery, unified namespace management, and portable agent identities represent the next frontier for infrastructure development.
@@
Implications for Trust-Critical Domains. Healthcare exemplifies why trust primitives matter: a patient care coordination agent must verify diagnostic agents’ credentials (authentication), ensure treatment recommendations haven’t been altered (integrity), and protect patient privacy during cross-hospital queries (confidentiality). NANDA Index’s W3C VC signatures and PrivateFactsURL support HIPAA-compliant audit trails [10] and privacy-preserving lookups. AGNTCY ADS’s Sigstore provenance enables ”only execute if verified” policies required by FDA software-as-medical-device regulations. Entra Agent ID’s Conditional Access integrates with existing healthcare IAM for zero-trust clinical workflows [1]. As AI agents expand into regulated sectors, registries become the root of trust for safe, accountable autonomous systems.
XI The Path Ahead: A Switchboard for the Agentic Web
@@
To move forward, our teams are deploying a working solution in two phases, inviting researchers and practitioners to join: Phase 1 establishes the NANDA Index at MIT as a neutral, consortium-governed switchboard for cross-organizational agent coordination; Phase 2 adds semantic discovery and federation capabilities through AGNTCY’s Agent Directory, enabling interoperability with heterogeneous registry types via DHT-based routing and OCI-aligned content addressing.